OpenStudy (danielbarriosr1):

The measure of one acute angle in a right triangle is four times the measure of the other acute angle. Write and solve a system of equations to find the measures of the acute angles.

OpenStudy (wolf1728):

The two acute angles must sum to 90° and one is 4 times larger than the other. So, angle 1 + 4*(angle 2) = 90° 5 * angle = 90° Angles = 18° and 72°

OpenStudy (danielbarriosr1):

But whats the equation? I mean.. i can see it but i dont know how to write it

OpenStudy (danielbarriosr1):

Please

OpenStudy (wolf1728):

Okay I guess I typed too quickly there. You have 2 angles that equal 90° One angle is 4 times larger than the other so angle + 4 * angle = 90° 5 * angle = 90° angle = 18° so other angle = 4*18° = 72°

OpenStudy (wolf1728):

Notice you do not need 2 equations to solve it.

OpenStudy (anonymous):

The measure of one acute angle in a right triangle is four times the measure of the other acute angle. Write and solve a system of equations to find the measures of the acute angles. First lest set up this equation The measure of one acute angle in a right triangle is four times the measure of the other acute angle x = one acute angle y = the other acute angle For the first equation, it tells us one acute angle is 4 time the other acute angle. We write this like x = 4y Now for the other equation and acute angle we need to know the total of all the angles and the total of all the angels on a right triangle is 180 degrees we set it up like so x + y + 90 = 180 A triangle has 3 sides but we do not know x or y but we know the 90 degrees angle. Now we substitute 4y for x because x = 4y Our equation now looks like 4y + y + 90 = 180 Now you can solve for y, which is one of the acute angles. Can you solve now or you need me to continue?

OpenStudy (danielbarriosr1):

Ok i got it thanks

OpenStudy (danielbarriosr1):

i dont know who to give the medal

OpenStudy (danielbarriosr1):

i mean i know how to do it now

OpenStudy (anonymous):

Once you solve for y you just put the answer of y into x = 4y So if y = 5 the other acute angle would be x = 4(5) , which x would = 20

OpenStudy (wolf1728):

MDoodler - solve it with just one equation

OpenStudy (danielbarriosr1):

i got y=18

OpenStudy (wolf1728):

That is the answer - look at both of my answers - you do not need 2 equations.

OpenStudy (danielbarriosr1):

ok

OpenStudy (wolf1728):

For the other angle if it is 4 times the other angle then it is 4 * 18 or ??

OpenStudy (danielbarriosr1):

72

OpenStudy (wolf1728):

Very good Daniel !

OpenStudy (danielbarriosr1):

Thanks. now i got it

OpenStudy (danielbarriosr1):

LOL

OpenStudy (anonymous):

And now that you got y = 18 all you had to do is insert y into x = 4y which would give you x, which is x = 4(18) = 72 Both ways work.

OpenStudy (danielbarriosr1):

Thanks

OpenStudy (anonymous):

Also @wolf1728 learn to explain and not just give answers like you did at first. It does not help them by just giving them the answers.

OpenStudy (danielbarriosr1):

Thats true

OpenStudy (danielbarriosr1):

But now i got it